# Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty

### v2

@@ I Introduction
    We provide both qualitative and quantitative insights into how positive and negative predictions influence the model’s reliability. Our analysis reveals that, unlike positives, negative predictions are often poorly calibrated and therefore should be handled separately. Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use (Section IV).
@@
    Identifying ground-truth positives and negatives ideally requires a matching process based on ground-truth annotations, which are unavailable at test time. Consequently, an alternative framework for identifying positive predictions is needed. However, we show that existing performance and calibration metrics—such as average precision [16, 17, 18] and expected calibration error [8, 19, 20]—are inadequate for this purpose. To address this, we introduce a new calibration metric, object-level calibration error (OCE), which assesses calibration error along ground-truth objects rather than predictions. We demonstrate that OCE is more suitable for both ranking different models by their calibration qualities and identifying the positive predictions (Section V).
  3. 3.
    Based on our findings, we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric. We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios (Section VI).
In addition to addressing the three primary research questions, we present a comparative analysis of standard post-processing methods for identifying positive predictions in DETR from an uncertainty quantification (UQ) perspective (Section VII). Our results indicate that thresholding on the predictive confidence score is more effective than widely used techniques such as top-$k$ selection or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
II Related Work
Calibration. Model calibration refers to how well a model’s predicted confidence scores align with the actual likelihood of those predictions being correct. In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy. For instance, if a model predicts an event with 80% confidence, it should be correct around 80% of the time for that event to be considered well-calibrated. Achieving good calibration is crucial because it improves trust in the model’s predictions, guides more informed decision-making, and enables better risk assessment in real-world applications.
To evaluate the alignment, they measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin. However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes. Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction. To address these challenges, D-ECE [8] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$). On the other hand, LaECE [21] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
However, as we demonstrate, DETR exhibits poor calibration on negative predictions, regardless of its actual reliability. In practice, users often focus only on positive samples. Yet, when model calibration is evaluated across the entire set of predictions, including the negatives, the resulting evaluation becomes highly biased and therefore unreliable. To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$. As detailed in later sections, a threshold of $0.3$ approximates the optimal positive predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable. However, the optimal threshold is not guaranteed to be always $0.3$ for other models like UP-DETR, which introduces potential risks. In response, our paper proposes OCE, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
Uncertainty Quantification. Several studies focus on out-of-distribution (OOD) identification in object detection models. For example, [22] propose a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions. [23] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data. Similarly, [19] employ an auxiliary detection model capable of expressing its confidence. Other works, including [24] and [25], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
To the best of our knowledge, the aforementioned existing UQ techniques primarily focus on prediction-level analysis. Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection models. In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs. Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
III Preliminaries
Figure 2: A diagram of the DETR architecture. An input image is first processed through a CNN backbone to generate a 2D feature representation. This representation is then passed to the Transformer encoder, which extracts feature vectors. These feature vectors are sent to the decoder, which receives $M$ learned object queries together. The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
III-A Detection Transformer (DETR)
We consider a test image $\bm{x}$ and denote $\mathcal{D}_{\bm{x}}$ as the set of ground truth objects present in the image. Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{D}}_{\theta}(\bm{x})$. Each prediction $d\in\hat{\mathcal{D}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions. In addition to the features extracted by the encoder, the decoder’s input consists of $M$ (typically several hundred) learnable embeddings, also known as object queries. Each decoder layer is composed of a self-attention module among object queries and a cross-attention module between each object query and the features. After processing the queries through several decoder layers, the model produces the $M$ final representation vectors that are converted into bounding boxes and class labels via a shared feedforward network, $f_{\phi}$. Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set. We refer to Figure 2 for an illustration.
It is noteworthy that the encoder follows the common structure of standard computer vision models, whose reliability has been relatively widely explored [26, 27, 28, 29]. This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [30]. However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction. Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of DETR.
III-B Bipartite Matching
Since the number of queries in DETR, $M$, is much higher than the number of annotated objects, DETR matches each object with the corresponding best model prediction during its training. To compute this optimal matching for the predictions in a given image, a bipartite matching algorithm is applied. More specifically, a matching cost between each pair of a given prediction and an object is defined as follows:
$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (1)
where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$. $\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [31] loss between bounding boxes. After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm [32, 7].
III-C Expected Calibration Error (ECE) Metrics
D-ECE. For a given object detection model, detection expected calibration error (D-ECE) [8] quantifies how closely the model’s predicted confidences align with its observed precision. Specifically, let $\hat{\mathcal{D}}$ be the set of all detections that the model produces on the validation set. Each detection has an associated predicted confidence score $\hat{p}\in[0,1]$. To compute D-ECE, the confidence space $[0,1]$ is partitioned into $J$ bins, for instance via uniform intervals $\big{[}0,\tfrac{1}{J}\big{)},\big{[}\tfrac{1}{J},\tfrac{2}{J}\big{)},\dots,\big{[}\tfrac{J-1}{J},1\big{]}$. Let $\widehat{\mathcal{D}}_{j}\subseteq\widehat{\mathcal{D}}$ be the subset of detections falling into the $j$-th bin. Denote by $\bar{p}_{j}$ the average confidence of detections in bin $j$, and by $\mathrm{precision}(j)$ the empirical precision in bin $j$. A detection is considered a true positive if it has the correct predicted class label and its IoU with the ground-truth box exceeds a given threshold $\tau$. The empirical precision $\mathrm{precision}(j)$ is thus the fraction of detections in $\widehat{\mathcal{D}}_{j}$ that are true positives.
D-ECE then aggregates the absolute difference between each bin’s average confidence and observed precision, weighted by bin size, and is defined as:

### v3

@@ IV-C Theoretical Analysis of the Optimal Prediction Strategy
    for the optimal positive prediction ($i=p$): $\hat{p}^{*}_{i}=p_{\text{gt}}$ (i.e., $\hat{p}^{*}_{i}(c)=p_{\text{gt}}(c)$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=0$) and the box $\hat{b}^{*}_{i}$ minimizes $\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b}_{i},b_{\text{gt}})]$, and
@@
    for any optimal negative prediction ($i\neq p$): $\hat{p}^{*}_{i}(c)=0$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=1$.
Proof.
See Section B-B. ∎
From Optimal Set to Strategy under Ambiguity. The previous propositions formally show that the Hungarian loss incentivizes a deterministic specialization. The optimal state is not a symmetric hedge where multiple queries compete, but for one query to emerge as a specialist that becomes a perfectly calibrated predictor of the object’s true class distribution ($p_{\text{gt}}$). All other queries are simultaneously optimized to predict the background class, minimizing their own loss to zero.
This finding implies that the model is incentivized to develop an internal strategy to achieve this state. Ideally, the model would learn to designate a specific object query (which we conceptually refer to as the primary query) to become the optimal positive prediction $p$, and the remaining ones (secondary queries) to become the optimal negatives. As we discussed in Section IV-B, this is architecturally feasible, as DETR can leverage its attention mechanism to orchestrate predictions: Throughout the decoder layers, the model can identify the most promising object query and progressively specialize it, while simultaneously suppressing the others.
However, in practice, a critical ambiguity arises. While the Hungarian matching algorithm $\hat{\sigma}$ is deterministic for a deterministic annotation, the ground truth itself is often ambiguous. Consequently, the assignment of queries to targets becomes a stochastic process. We formalize this by defining $P_{i}$ as the probability that the $i$-th query is matched to the ground-truth object:
$P_{i}\triangleq\mathbb{P}_{y_{\text{gt}}\sim Y}(\hat{\sigma}(1)=i).$ (5)
In this setting, the extreme specialization is now suboptimal because it incurs an infinite loss if the primary prediction is matched to the background. Therefore, to minimize its expected loss under this uncertainty, each query must find an optimal prediction $\hat{p}_{i}^{*}$ that “hedges” against this matching uncertainty. The expected loss $J_{i}$ for query $i$ is decomposed as:
$J_{i}$ $=\underbrace{P_{i}\cdot\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b}_{i},b_{\text{gt}})\mid\hat{\sigma}(1)=i]}_{\text{Expected Box Loss Term}}$
$\quad\underbrace{-P_{i}\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{i}(c))-(1-P_{i})w_{\emptyset}\log(\hat{p}_{i}(\emptyset))}_{\text{Expected Classification Objective }J(\hat{p}_{i})}.$
Minimizing $J_{i}$ with respect to $\hat{p}_{i}$ is equivalent to minimizing the expected classification objective $J(\hat{p}_{i})$. Proposition 3 characterizes the optimal strategy to minimize the objective.
Proposition 3 (Optimal Strategy under Uncertainty).
For a query with a matching probability $P_{i}$ to the ground-truth object whose class distribution is $p_{\text{gt}}$, the class prediction that minimizes the expected loss is $\hat{p}_{i}^{*}$, where
  1. 1.
    the prediction for any foreground class $c\in\mathcal{C}$ is:
    $\hat{p}_{i}^{*}(c)=p_{\text{fg}}^{*}(P_{i})p_{\text{gt}}(c).$ (6)
  2. 2.
    the prediction for the background class is
    $\hat{p}_{i}^{*}(\emptyset)=1-p_{\text{fg}}^{*}(P_{i}),$
where $p_{\text{fg}}^{*}(P_{i})$ is the optimal foreground probability, defined as
$p_{\text{fg}}^{*}(P_{i})=\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}.$ (7)
Proof.
See Section B-C. ∎
Based on the result from Proposition 3, we analyze two characteristics in their prediction strategy:
  • Perfect Specialization: Consider the scenario where the model is highly certain about the image context and thus about which query will be matched. In this scenario, the matching probability for the primary—which will eventually become the optimal positive—query index is perceived as 1 (i.e., $P_{p}\to 1$), while it vanishes for all others (i.e., $P_{i}\to 0$ for $i\neq p$). In this scenario, the strategy recovers the optimal prediction set (Proposition 2), where solely the $p$-th query becomes a perfectly calibrated predictor of the object ($p_{\text{gt}}$).
  • Hedging: When the model is uncertain about the matching query (i.e., $P_{i}$ is dispersed), it optimally “hedges.” The foreground probability for the primary query is suppressed ($p_{\text{fg}}^{*}(P_{p})<1$), balancing the risk between foreground and background assignments.
In practice, matching probabilities are skewed: $P_{p}$ is close to 1 for the primary query, and $P_{i}$ is close to 0 for secondary queries: $i\neq p$. Thus, DETRs prioritize specialization while retaining a hedging mechanism proportional to their uncertainty. We term this the Specialist Strategy. This jives with the empirical trends observed in Figures 3 and 4.
Figure 5: Empirical cumulative distribution function (ECDF) plots for maximum confidence score over foreground classes and maximum IoU over ground truth objects, comparing optimal positive predictions (blue) and optimal negative predictions (red), in UP-DETR and Cal-DETR on the COCO dataset. (a, b) Confidence: The optimal negative prediction confidence is overwhelmingly clustered at zero. Optimal positive predictions span the full confidence range, acting as the system’s calibrated predictors. (c, d) Localization: Optimal positive predictions are highly accurate (high IoU), which is the necessary condition for the model to designate a primary query. Optimal negative predictions maintain decent localization, which is not a failure but the optimal risk-averse strategy.
IV-D Numerical Analysis
Setup. To provide quantitative support for our claim, we conducted experiments using four DETR variants: UP-DETR [35], Deformable-DETR (D-DETR) [14], Cal-DETR [12], and DINO [15]. We also investigate with Faster R-CNN [2] and YOLOv3 [redmon2018yolov3, 3]. Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set.
Empirical Analysis of Prediction-Level Distributions. We now empirically analyze the distributions of both confidence and localization accuracy to validate our theoretical findings. Ideally, this analysis would compare the primary prediction (the query intended to be the optimal positive) against secondary predictions. However, since the model’s internal designation is latent and not directly observable, we utilize the outcomes of the Hungarian matching as effective proxies. Specifically, we analyze optimal positive predictions (those matched to the ground truth) as the observable realization of primary prediction, and optimal negative predictions (those matched to the background) as the proxy for secondary predictions.
First, we investigate the confidence scores. We plot the empirical cumulative distribution function (ECDF) of the maximum foreground probability ($\max_{c\neq\emptyset}\hat{p}_{i}({c})$) in Figures 5a and 5b. This indicates that the optimal negative predictions’ confidence scores are heavily concentrated in the near-zero regime, effectively being suppressed, in contrast to the widely dispersed confidence distribution observed for optimal positive predictions.
Next, we perform a parallel analysis on localization accuracy in Figures 5c and 5d. More specifically, for each prediction, we measure the maximum IoU with respect to the ground truth objects in the image. The results for optimal positive predictions confirm that the vast majority exhibit high accuracy. On the other hand, the analysis of optimal negative predictions reveals a critical distinction: despite their suppressed confidence, they frequently maintain substantial localization accuracy. It reflects a risk-averse strategy where the model maintains precise backup bounding boxes to mitigate the high penalty of potential matching errors, ensuring training stability. As shown in Figure 13, these observed trends are consistent across various DETR variants and other detectors.
We remark that this is not a model failure, but rather a risk-averse strategy induced by the loss landscape. In a stochastic setting where $y_{\text{gt}}$ varies, there is always a non-zero probability that secondary queries get matched to the ground-truth object. If the secondary query predicts a random box, its matching to the foreground object elevates the expected box loss. Thus, to minimize the expected total loss, the model is incentivized to set secondary queries to predict a decent bounding box rather than a random box.
Figure 6: Flowchart of our evaluation framework. Two components vary across configurations: (1) the model ($\theta$), which generates a full set of predictions ($\hat{\mathcal{Y}}_{\theta}(x)$), and (2) the post-processing algorithm ($\mathcal{A}$), which selects a positive set ($\hat{\mathcal{S}}_{\theta}(x)$) from these predictions. In addition, we can obtain the optimal positive sets ($\mathcal{S}_{\theta}(x)$) by passing the full prediction set ($\hat{\mathcal{Y}}_{\theta}(x)$) and ground-truth object annotations ($\mathcal{Y}$) to the Hungarian Algorithm.
Figure 7: Flowchart of our method for evaluating calibration error. A calibration metric ($\mathcal{M}$) compares either given positive sets ($\hat{\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}})$) or optimal positive sets (${\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}})$) with ground-truth object annotations ($\mathcal{Y}_{\text{test}}$) to get calibration error ($\mathcal{M}(\hat{\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}}),\mathcal{Y}_{\text{test}})$) or optimal calibration error ($\mathcal{M}({\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}}),\mathcal{Y}_{\text{test}})$), respectively.
V Quantifying Calibration in DETR
V-A Implications of the Specialist Strategy for Evaluation
Our analysis reveals that DETR’s training objective incentivizes a specialist strategy. This strategy assigns two fundamentally different roles to the predictions:
  1. 1.
    A small subset of primary predictions (intended to be matched with foreground objects) is trained to be calibrated predictors, matching the true class distribution $p_{G}$.
  2. 2.
    The vast majority of secondary predictions are trained to be background classifiers, intentionally suppressing their foreground confidence to near-zero.
This specialist strategy introduces a fundamental ambiguity in downstream decision-making. When a prediction exhibits low confidence, it becomes indistinguishable: is it a positive prediction correctly expressing high uncertainty about an ambiguous object, or is it a secondary prediction whose confidence has been artificially suppressed? This distinction is paramount, particularly in safety-critical domains.
Safety-Critical Implication For an autonomous vehicle, a low-confidence “person” prediction could be a truly ambiguous object (like a person depicted on a billboard or on a vehicle), which the model is correctly uncertain about. However, it could also be a definite pedestrian identified by a secondary prediction, whose confidence was artificially suppressed to near-zero as part of the model’s behavior, resulting from optimizing the Hungarian loss (as discussed in Section IV). In the latter case, relying on this suppressed score could lead the system to catastrophically misjudge the risk posed by a real person.
Ideally, at test time, we aim to filter out the secondary predictions and retain only the primary ones. However, identifying the primary predictions is non-trivial because the model’s internal designation is latent. Consequently, practitioners rely on a post-processing algorithm (e.g., confidence thresholding) to select a subset of predictions serving as an estimate of the primary set.
While one might consider evaluating calibration directly on the optimal positive predictions, identifying this set requires ground-truth labels, making it infeasible at test time. Evaluating against this oracle set is useful for analysis, but it does not perfectly reflect the actual deployment conditions, where the system must rely solely on a post-processed subset.
These issues create a critical challenge for practical deployment: we must evaluate calibration quality using the same post-processing algorithm that will be applied at test time. Therefore, practitioners should assess the entire (model + post-processing) pipeline jointly to identify configurations that produce reliable positive sets, as illustrated in Figures 6 and 7, rather than evaluating the model in isolation. To address this, we propose a new calibration metric that takes as input any positive set produced by a post-processing algorithm, enabling such joint evaluation.
V-B Proposed Metric: Object-Level Calibration Error
Notation. For each image $x_{l}$ having $N_{l}$ objects, we consider a set of ground-truth object annotations $\mathcal{Y}_{l}=\{y_{l,j}=(c_{l,j},b_{l,j})\}_{j=1}^{N_{l}}$, and a set of DETR predictions: $\hat{\mathcal{Y}}_{\theta}(x_{l})=\{\hat{y}_{l,i}=(\hat{p}_{l,i},\hat{b}_{l,i})\}_{i=1}^{M}$ where $M$ represents the number of object queries. Let $\mathcal{Y}_{\text{test}}=\big(\mathcal{Y}_{l}\big)_{l=1}^{N_{\text{test}}}$ denote the set of all ground-truth objects across all $N_{\text{test}}$ test images $\mathcal{X}_{\text{test}}=\big(x_{l}\big)_{l=1}^{N_{\text{test}}}$.
Definition 1.
Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(x_{l})\subseteq\hat{\mathcal{Y}}_{\theta}(x_{l})$, that is generated by post-processing algorithm $\mathcal{A}$ from the full prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{A}\circ\hat{\mathcal{Y}}_{\theta}$. We define an object-level calibration error (OCE) as the average Brier score per object:
$\mathrm{OCE}_{\tau}\big(\hat{\mathcal{S}}_{\theta},\mathcal{Y}_{\text{test}}\big)=\frac{1}{|\mathcal{Y}_{\text{test}}|}\sum_{y_{l,j}\in\mathcal{Y}_{\text{test}}}\mathrm{Brier}_{\tau}\big(\hat{\mathcal{S}}_{\theta}(x_{l}),y_{l,j}\big),$ (8)
$\mathrm{Brier}_{\tau}\big(\hat{\mathcal{S}}_{\theta}(x_{l}),y_{l,j}\big)=\sum_{c=1}^{C}\big(\mathbbm{1}(c=c_{l,j})-\bar{p}^{(\tau)}_{l,j}(c)\big)^{2},$
$\bar{p}^{(\tau)}_{l,j}(c)=\frac{1}{|\mathcal{Q}^{(\tau)}_{l,j}|}\sum_{q\in\mathcal{Q}^{(\tau)}_{l,j}}\hat{p}_{l,i}(c),$
where $\mathcal{Q}^{(\tau)}_{l,j}$ is a set of prediction indices that has bounding box overlap with the ground-truth object $y_{l,j}$:
$\mathcal{Q}^{(\tau)}_{l,j}$ $\triangleq\{i\mid\text{IoU}(b_{l,j},\hat{b}_{l,i})\geq\tau\}.$
When $|\mathcal{Q}^{(\tau)}_{l,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$. Following [18, 20], we use IoU thresholds of $\tau=0.5,0.75$ and report the average score as $\mathrm{OCE}(\cdot)=\mathrm{OCE}_{0.5}(\cdot)+\mathrm{OCE}_{0.75}(\cdot)$.
The introduced object-level calibration error (OCE) has two desirable characteristics derived from its object-centric design. Unlike prediction-level metrics (e.g., D-ECE) that average errors over the selected predictions, OCE averages errors over the ground-truth objects. Consequently, the metric effectively penalizes two critical failure modes in recovering the primary predictions. (1) Retaining secondary predictions: OCE penalizes subsets containing secondary predictions—those that maintain systematically low confidence despite accurate localization—as they fail to contribute correctly to the assigned objects. (2) Missing primary predictions: OCE penalizes subsets that are too aggressive and discard primary predictions, leaving ground-truth objects unmatched and incurring a maximum penalty. Thus, unlike existing prediction-level ECE metrics, OCE effectively measures how well the post-processed subset approximates the set of reliable predictions.
Remark (complexity). Computing OCE only requires IoU checks between GT boxes and predicted boxes; precomputing an $N_{i}\times M$ IoU matrix per image makes the additional cost negligible relative to standard evaluation.
V-C Numerical Analysis
Table 1: The Pearson correlation coefficient between the proposed OCE and established calibration metrics, D-ECE and LA-ECE0, on the optimal positive predictions () across six models—UP-DETR, D-DETR, Cal-DETR, DINO, Faster R-CNN, and YOLOv3—and three data distributions. The consistently very strong correlations confirm that OCE effectively captures the same underlying concept of calibration quality as D-ECE and LA-ECE0.
Predictions Metrics COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average $\pm$ Std.